OpenStudy (anonymous):

No idea how to solve this.. The half-life of radium is 1690 years. If 10 grams is present now, how much will be present in 50 years?

OpenStudy (anonymous):

here first of all find the disintegrating constant

OpenStudy (anonymous):

i am not sure abt the formula but it is somewhat like disintegrating constant =0.693/(hal-life time)

OpenStudy (anonymous):

then use ln (amt present after t yrs/amt initially present) = -(disintegrating constant)* time

OpenStudy (anonymous):

@Kira_Yamato Can you explain how you did it?

OpenStudy (anonymous):

radio active decay problems modelled by : \(\large y=Ce^{kt} \) where C is your initial amount (10 grams), and k is your "disintegrating" constant as explained by @matricked. you'll need to find k so you'll need to solve the equation: \(\large 5=10e^{k \cdot 1690} \) because the half-life is 1690 years, there will only be 5 grams left of the original 10 grams when 1690 years have elapsed. This is what @Kira Yamato did above. k=-(ln2)/1690. So the model for the half-life decay is: \(\large y=10e^{(-\frac{ln2}{1690}\cdot t)} \) To answer your question, use a calculator to find y when t=50 years.

OpenStudy (anonymous):

@dpaInc @toadytica305 @Kira_Yamato that's how you solve the problem. we do not need to stick a negative sign to K because when you solve K will be negative.